Mr. Alvarez, Mr. Brown, and Mr. Chen are partners in a buisness that generated $500,000 in profits. You know Mr. Alvarez owns 40% of the buisness and that Mr. Chen has received 175,000 of the profits.
Find the amount of the profits that Mr. Alvarez and Mr. Brown should receive

Answer by Theo(13342)

profits = 500,000
alvarez owns 40% of the business so he should get 40% of the profits.
.4 * 500,000 = 200,000

chen gets 175,000 of the profits.

brown gets the rest.
500,000 - 200,000 - 175,000 = 500,000 - 375,000 = 125,000

alvarez gets 200,000
chen gets 175,000
brown gets 125,000

alvarez owns 40% of the business
chen must own 35% of the business
brown must own 25% of the business

This would be true, if they split the profits according to their percent ownership in the business.

That's the assumption I made.